LA. and Paula file as married taxpayers. In August of this year, they received a 5,580 refund of state income taxes that they paid last year.
How much of refund if any, must L. A and Paula included in gross income under the following independent scenario? Assume the standard deduction last year was 12,700.
Last year L. A. and Paula had itemized deductions of 10100,and they chose to claim the standard deduction.
How much refund to be included
Answer #1

Rate of interest = 15%

15 / 100 x ? = 7,600

15 x ? = 7,600 x 100

15 x ? = 760,000

? = 760,000 / 15 = 50,667

Total Income balancing figure

$76,367.00

- Adjustments

$0.00

- Deductions

$25,700.00

- Exemptions

$

= Taxable Income

$50,667.00

Tax

$7,600.00

+ Alternative Minimum Tax

$0.00

- Credits

$0.00

+ Other Taxes

$0.00

- Withholdings & Estimated Payments

$0.00

- Refundable Credits

$0.00

= Refund

$6,920.00